# Multi-Hypothesis Prediction for Portfolio Optimization: A Structured Ensemble Learning Approach to Risk Diversification

### v2

@@ 3.3 Multi-Hypothesis Portfolio Optimization with Structured Radial Basis Function Network (s-RBFN)
The structured data set is used as input for a radial basis function network, with each $j^{th}$ predictor or hypothesis $f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})=\boldsymbol{R}_{j}(\varepsilon)$ associated with a particular basis function $\phi\left(\boldsymbol{R}_{j}(\varepsilon),\mu_{j},\sigma_{j}\right)$, that is, a map $\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right):\mathbb{R}^{N\times M}\to\mathbb{R}^{N\times M}$ is obtained by applying the basis function $\phi\left(\cdot\right)$ to each element of $\boldsymbol{R}(\varepsilon)$. For example, the Gaussian basis function $\phi\left(\cdot\right)=\exp{\left(\frac{-1}{2\sigma_{j}^{2}}\left|{R}_{ij}(\varepsilon)-\mu_{j}\right|^{2}\right)}$ has been used where the parameters of the centers $\mu_{j}$ and the scales $\sigma_{j}$ for the basis functions are calculated from each column of $\boldsymbol{R}(\varepsilon)$ as the mean and standard deviation, respectively. The s-RBFN formulation can now be expressed in matrix form as follows:
$\boldsymbol{\hat{R}}_{t}=\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)\boldsymbol{w}=$
@@
$=\left[\begin{matrix}\phi\left({R}_{11}(\varepsilon),\mu_{1},\sigma_{1}\right)&\ldots&\phi\left({R}_{1M}(\varepsilon),\mu_{M},\sigma_{M}\right)\\ \vdots&\ddots&\vdots\\ \phi\left({R}_{N1}(\varepsilon),\mu_{1},\sigma_{1}\right)&\ldots&\phi\left({R}_{NM}(\varepsilon),\mu_{M},\sigma_{M}\right)\\ \end{matrix}\right]\left[\begin{matrix}w_{1}\\ \vdots\\ w_{M}\\ \end{matrix}\right]$ (12)
@@
The optimal s-RBFN parameters are obtained by least-squares with the regularization parameter $\lambda_{s}$ and target portfolio $\boldsymbol{R}_{t}$:
$\boldsymbol{w}=\left(\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)^{\rm T}\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)+\lambda_{s}\ast\boldsymbol{I}_{(mxm)}\right)^{-1}\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)^{\rm T}\boldsymbol{R}_{t}$ (13)
@@
The portfolio weights are the scaled s-RBFN parameters, which can be scaled in various ways, including ensuring they sum to one and remain within the range of 0 to 1. This particular scaling reflects the constraints of long-only portfolios, satisfying $\mathbf{1}^{\top}\mathbf{w}=1$ and $0\leq\mathbf{w}\leq 1$.
The time series returns of the target portfolio used to train the structured portfolio ensemble and compute the portfolio weights have the expression $\boldsymbol{R}_{t}=\boldsymbol{w}^{T}\boldsymbol{R}(\varepsilon)$. If (13) is reduced to:
$\boldsymbol{w}=\boldsymbol{P}\boldsymbol{R}_{t}=\boldsymbol{P}\boldsymbol{w}^{\top}\boldsymbol{x}$ (14)
@@
with $\boldsymbol{P}=\left(\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)^{\top}\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)+\lambda_{s}\ast\boldsymbol{I}_{(m\times m)}\right)^{-1}\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)^{\top}$. The target portfolio time-series returns in the training set in matrix form is given by:
$\boldsymbol{R}_{t}=\boldsymbol{w}_{\text{opt}}^{\top}\boldsymbol{x}$ (15)
with $\boldsymbol{w}_{\text{opt}}$ being the optimal weights and $\boldsymbol{x}$ the time series of the portfolio constituents. For $\boldsymbol{w}$ to be optimal in equation (13), $\boldsymbol{R}_{t}$ must be given by (15), yielding equation (14).
@@
To ensure that the optimization is based on the shape of the constituents’ hypothesis space and the degree of diversification in the structured data set of predictions that have already been incorporated into $\boldsymbol{P}$, a canonical basis of hypotheses, each of which corresponds to one of the portfolio constituents, is selected. For that, the weights in equation (14) are selected such that the portfolio $\boldsymbol{R}_{t}$ is equal-weighted, $\boldsymbol{w}=\left\{\frac{1}{\|\boldsymbol{x}\|}\right\}$. This ensures that the optimization and final solution for the portfolio weights are based on the information in $\boldsymbol{P}$ and not other pre-specified targets:
$\boldsymbol{w}=\boldsymbol{P}\frac{\boldsymbol{x}}{\|\boldsymbol{x}\|}$ (16)
The Centroid Combiner Rule, adapted to the problem of portfolio optimization, can be derived from the Bias-Variance-Diversity decomposition in (10) (Wood et al., 2024):
$\boldsymbol{\hat{R}}_{t}=\arg\min\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{R}_{t},\boldsymbol{R}_{j}),$
which is the equal-weighted portfolio, described in the literature as the 1/N rule, and represents the naive portfolio diversification rule (DeMiguel et al., 2009).
@@
In the mean-variance framework, the key plug-in component is the covariance matrix, which is based exclusively on the input data set $\boldsymbol{x}$ and does not provide a guarantee of achieving portfolio diversification outside the sample. In contrast, the PSEM framework incorporates portfolio diversification directly into the modeling process by treating assets as hypotheses and the portfolio as an ensemble predictor, leveraging the diversity among individual predictors during training. The structured data set $\boldsymbol{R}(\varepsilon)$, constructed from predictions generated by models using the same input data as the mean-variance model, enables the use of a continuous diversification parameter $10<\varepsilon<$ 1. This allows for the creation of multiple diversified plug-in data sets from the same input and for portfolio diversification to be adjusted prior to weight optimization.
@@
Both approaches involve matrix multiplication as a core computational operation in portfolio optimization, rendering them computationally equivalent in this respect. In the PSEM framework using the s-RBFN model, the input data $\boldsymbol{x}$ is replaced with the diversified structured data set of predictions $\boldsymbol{R}(\varepsilon)$, which is augmented through the use of basis functions. These basis functions—such as radial, Gaussian, or other types—introduce additional flexibility to the s-RBFN model, enabling it to capture more complex non-linear relationships within the data.
@@
Figure 2: Systems representation comparison of the Mean-Variance approach to portfolio optimization from Modern Portfolio Theory (Markowitz, 1952) (Left) versus the s-RBFN optimized by least squares (Dominguez et al., 2025) (Right). The portfolio constituents’ returns, denoted as $\boldsymbol{x}=\{\boldsymbol{x_{j}}\}_{j=1}^{M}$, and the target portfolio returns, denoted as $\boldsymbol{y}$ in the left case are assumed to be normally distributed. The target portfolio returns in the right case are $\boldsymbol{R}_{t}$. The remaining parameters and variables include the portfolio weights $\boldsymbol{w}$, the diversity parameter $10<\varepsilon<$ 1 used for the structured data set $\boldsymbol{R}(\varepsilon)$ formation, the individual predictors $\{f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x_{j}})\}_{j=1}^{M}$, parameterized by $\Theta=\{\boldsymbol{\theta}_{j}\}_{j=1}^{M}$, and the s-RBFN model basis functions $\boldsymbol{\Phi}$.
@@
The mean-variance approach is a fully supervised method, where the portfolio optimization process is guided by a predefined target return series. In contrast, the PSEM framework operates within a semi-supervised learning paradigm, as it does not rely on a predefined target. Instead, the optimal portfolio weights are determined based on the hypotheses generated by the predictors and their mixtures during training. Consequently, to provide a canonical reference for the hypothesis space, it is necessary to employ an equal-weighted portfolio as a target portfolio ($\boldsymbol{R}_{t}=\frac{1}{||\boldsymbol{x}||}\boldsymbol{x}$ in Figure 2). This ensures that the decision-making is governed solely by the structured data set $\boldsymbol{R}(\varepsilon)$ and the s-RBFN model, thereby avoiding any dependence on a specified target.
@@ 4.2.2 Including Diversity in the Asset or Hypothesis Selection Stage
In this section, the same experiments are conducted, but here the $M$ portfolio constituents or hypotheses are randomly selected from the top $m=M\times\gamma$ stocks based on the 1-month cumulative forecast ranking, where $\gamma$, the asset selection diversity parameter, is a multiplier that scales the candidate pool relative to the final selection size $M$. This approach seeks to test the hypothesis that out-of-sample diversification in portfolio allocation is achieved not only during the optimization process—whether through the inclusion of the parameter $\varepsilon$ in this structured prediction model framework or through other plug-in methods without such aid—but also at the stage of hypothesis selection, when choosing assets or stocks prior to optimizing portfolio weights. For that, the experiments are repeated for different values of the multiplier, including the no-diversity case, which coincides with the results in the previous section, with $\gamma=1,2,3,5$ for both one-step-ahead and multi-step-ahead cases, using the s-RBFN model with radial basis functions.
@@ 5 Conclusion and Future Work
(Markowitz, 1952) that are the basis of modern finance can be understood through structured models for multiple hypotheses prediction, providing new insights. By framing the problem this way, the diversity of individual predictors can be linked with portfolio diversification, offering a fresh perspective based on learning in multiple hypothesis prediction settings. The proposed PSEM strategy enables a structured framework for portfolio optimization, using as portfolio target the arithmetic combiner or equal-weighted portfolio. To the authors’ knowledge, the first plug-in method that parametrically controls out-of-sample diversification prior to the decision-making is introduced. This is useful for practitioners as they can access many diversified versions of the structured data set with different levels of diversification before the portfolio optimization.

### v3

@@ 3.3 Multi-Hypothesis Portfolio Optimization with Structured Radial Basis Function Network (s-RBFN)
In this case, the structured dataset is used as input to a radial basis function network (RBFN), where each $j$-th predictor $f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})$ is associated with a specific basis function $\phi_{j}(\hat{x}_{ij}(\varepsilon),\mu_{j},\sigma_{j})$. The basis function $\phi_{j}$ is applied elementwise to the entries of the $j$-th input vector of predictions $\boldsymbol{\hat{x}}_{j}(\varepsilon)\in\mathbb{R}^{N}$, with $\mu_{j}$ and $\sigma_{j}$ representing the center and scale parameters specific to the $j$-th predictor. This defines a transformation map $\boldsymbol{\Phi}(\boldsymbol{\hat{x}}(\varepsilon)):\mathbb{R}^{N\times M}\to\mathbb{R}^{N\times M}$, where $\boldsymbol{\Phi}(\boldsymbol{\hat{x}}(\varepsilon))_{ij}=\phi_{j}(\hat{x}_{ij}(\varepsilon),\mu_{j},\sigma_{j})$. For example, the Gaussian basis function $\phi_{j}\left(\cdot\right)=\exp{\left(\frac{-1}{2\sigma_{j}^{2}}\left|{\hat{x}}_{ij}(\varepsilon)-\mu_{j}\right|^{2}\right)}$ can be used where the parameters of the centers $\mu_{j}$ and the scales $\sigma_{j}$ for the basis functions are calculated from each column of $\boldsymbol{\hat{x}}(\varepsilon)$ as the mean and standard deviation, respectively. The s-RBFN formulation can now be expressed in matrix form as follows:
@@
$\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)^{T}\boldsymbol{w}=\left[\begin{matrix}\phi\left(\boldsymbol{\hat{x}}_{11}(\varepsilon),\mu_{1},\sigma_{1}\right)&\ldots&\phi\left(\boldsymbol{\hat{x}}_{1M}(\varepsilon),\mu_{M},\sigma_{M}\right)\\ \vdots&\ddots&\vdots\\ \phi\left(\boldsymbol{\hat{x}}_{N1}(\varepsilon),\mu_{1},\sigma_{1}\right)&\ldots&\phi\left(\boldsymbol{\hat{x}}_{NM}(\varepsilon),\mu_{M},\sigma_{M}\right)\\ \end{matrix}\right]\left[\begin{matrix}w_{1}\\ \vdots\\ w_{M}\\ \end{matrix}\right]$ (5)
@@
The optimal s-RBFN parameters are obtained by least-squares with the regularization parameter $\lambda_{s}$ and target portfolio $\boldsymbol{r}$:
$\boldsymbol{w}=\left(\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)^{\rm T}\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)+\lambda_{s}\ast\boldsymbol{I}_{(mxm)}\right)^{-1}\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)^{\rm T}\boldsymbol{r}$ (6)
@@
The portfolio weights are the scaled s-RBFN parameters, which can be scaled in various ways, including ensuring they sum to one and remain each within the range of 0 to 1. This particular scaling reflects the constraints of long-only portfolios, satisfying $\mathbf{1}^{\top}\boldsymbol{w}=1$ and $0\leq\boldsymbol{w}\leq 1$. If (6) is reduced to:
$\boldsymbol{w}=\boldsymbol{P}\boldsymbol{r}$ (7)
@@
with $\boldsymbol{P}=\left(\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)^{\top}\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)+\lambda_{s}\ast\boldsymbol{I}_{(m\times m)}\right)^{-1}\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)^{\top}$. The target portfolio time-series returns is given by:
$\boldsymbol{r}=\boldsymbol{w}_{\text{opt}}^{\top}\boldsymbol{x}$ (8)
@@
with $\boldsymbol{w}_{\text{opt}}$ being the optimal weights and $\boldsymbol{x}$ the time series of the portfolio constituents. For $\boldsymbol{w}$ to be optimal in equation (6), $\boldsymbol{r}$ must be given by (8). To ensure in the PSEM that the optimization is guided by the structure of the constituents’ hypothesis prediction space and the degree of diversification within the structured dataset of predictions already incorporated into $\boldsymbol{P}$, a canonical basis of hypotheses is selected. For that, the weights in equation (7) are selected such that the target portfolio $\boldsymbol{r}$ is equal-weighted, $\boldsymbol{w}=\left\{\frac{1}{\|\boldsymbol{x}\|}\right\}$. This ensures that the optimization and final solution for the portfolio weights are based on the information in $\boldsymbol{P}$ and not on other pre-specified targets:
$\boldsymbol{w}_{\text{opt}}=\boldsymbol{P}\frac{\boldsymbol{x}}{\|\boldsymbol{x}\|}$ (9)
with $\frac{\boldsymbol{x}}{\|\boldsymbol{x}\|}$ forming the canonical basis of hypotheses. Additionally, the centroid combiner rule in the bias-variance-diversity decomposition for the case of the MSE loss in (2), when the ensemble is a weighted combination as in the s-RBFN case, is given by $\boldsymbol{\bar{r}}=\frac{1}{M}\sum_{j=1}^{M}w_{j}\phi(\boldsymbol{\hat{x}}_{j}(\varepsilon))$, which is equivalent to the equal-weighted portfolio $\frac{1}{M}\sum_{j=1}^{M}\boldsymbol{x}_{j}$, if the assets are taken as the hypotheses instead of the weighted basis functions.
@@
In the mean-variance framework, the key plug-in component is the covariance matrix, which is based exclusively on the input data set $\boldsymbol{x}$ and does not provide a guarantee of achieving portfolio diversification outside the sample. In contrast, the PSEM framework incorporates portfolio diversification directly into the modeling process by treating assets as hypotheses and the portfolio as an ensemble predictor, leveraging the diversity among individual predictors during training. The structured data set $\boldsymbol{\hat{x}}(\varepsilon)$, constructed from predictions generated by models using the same input data as the mean-variance model, enables the use of a continuous diversification parameter $10<\varepsilon<$ 1. This allows for the creation of multiple diversified plug-in data sets from the same input and for portfolio diversification to be adjusted prior to weight optimization.
@@
Both approaches involve matrix multiplication as a core computational operation in portfolio optimization, rendering them computationally equivalent in this respect. In the PSEM framework using the s-RBFN model, the input data $\boldsymbol{x}$ is replaced with the diversified structured data set of predictions $\boldsymbol{\hat{x}}(\varepsilon)$, which is augmented through the use of basis functions. These basis functions—such as radial, Gaussian, or other types—introduce additional flexibility to the s-RBFN model, enabling it to capture more complex non-linear relationships within the data.
@@
Figure 2: Systems representation comparison of the Mean-Variance approach to portfolio optimization from Modern Portfolio Theory (Markowitz, 1952) (Left) versus the s-RBFN optimized by least squares (Dominguez et al., 2025) (Right). The portfolio constituents’ returns, denoted as $\boldsymbol{x}$, and the target portfolio returns, denoted as $\boldsymbol{r}$ in the left case are assumed to be normally distributed. The remaining parameters and variables include the portfolio weights $\boldsymbol{w}$, the diversity parameter $10<\varepsilon<$ 1 used for the structured data set $\boldsymbol{\hat{x}}(\varepsilon)$ formation, the individual predictors $\{f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})\}_{j=1}^{M}$, and the s-RBFN model basis functions, $\boldsymbol{\phi}$ and $\boldsymbol{\Phi}$.
@@
The mean-variance approach is a fully supervised method, where the portfolio optimization process is guided by a predefined target return series. In contrast, the PSEM framework operates within a semi-supervised learning paradigm, as it does not rely on a predefined target. Instead, the optimal portfolio weights are determined based on the hypotheses generated by the predictors and their mixtures during training. Consequently, to provide a canonical reference for the hypothesis space, it is necessary to employ an equal-weighted portfolio as a target portfolio ($\boldsymbol{r}=\frac{1}{||\boldsymbol{x}||}\boldsymbol{x}$ in Figure 2). This ensures that the decision-making is governed solely by the structured data set $\boldsymbol{\hat{x}}(\varepsilon)$ and the s-RBFN model, thereby avoiding any dependence on a specified target.
@@ 4.2.2 Including Diversity in the Asset or Hypothesis Selection Stage
In this section, the same experiments are performed, but here the $M$ portfolio constituents or hypotheses are randomly selected from the top $m=M\times\gamma$ stocks based on the 1-month cumulative forecast ranking, where $\gamma$, the parameter of asset selection diversity, is a multiplier that scales the candidate pool relative to the final selection size $M$. This approach seeks to test the hypothesis that out-of-sample diversification in portfolio allocation is achieved not only during the optimization process—whether through the inclusion of the parameter $\varepsilon$ in this structured prediction model framework or through other plug-in methods without such aid—but also at the stage of hypothesis selection, when choosing assets or stocks prior to optimizing portfolio weights. For that, the experiments are repeated for different values of the multiplier, including the no-diversity case, which coincides with the results in the previous section, with $\gamma=1,2,3,5$ for both one-step-ahead and multi-step-ahead cases, using the s-RBFN model with radial basis functions.
@@ 5 Conclusion and Future Work
(Markowitz, 1952) that are the basis of modern finance can be understood through structured models for multiple hypotheses prediction, providing new insights. By framing the problem this way, the diversity of individual predictors can be linked with portfolio diversification, offering a fresh perspective based on learning in multiple hypothesis prediction settings. The proposed PSEM strategy enables a structured framework for portfolio optimization, using as portfolio target the arithmetic combiner or equal-weighted portfolio. To the authors’ knowledge, the first plug-in method that parametrically controls out-of-sample diversification prior to the decision-making is introduced. This is useful for practitioners, as they can access many diversified versions of the structured data set with different levels of diversification before portfolio optimization.